OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z + 4) = 125 ?

OpenStudy (anonymous):

Try taking natural log for both sides

OpenStudy (anonymous):

5^2(z+4)=5^3 by comparing the power 2(z+4)=3 2z+8=3 2z=3-8 2z=-5 z=-5/2

OpenStudy (anonymous):

if you need to find "z"

OpenStudy (anonymous):

Yes it is. Thank you very much